OpenStudy (anonymous):

Can someone please re-teach me??? Solve the equation. -5y-9=-(y-1)

OpenStudy (studygurl14):

First distribute the - to the y and -1

OpenStudy (anonymous):

So now its -5y-9=y+1?

OpenStudy (studygurl14):

Not quite. Don't forget to make the y negative.

OpenStudy (anonymous):

oh ok. so -5y-9=-y+1

OpenStudy (anonymous):

What next?

OpenStudy (studygurl14):

Now, combine like terms. \(\large\color{blue}{-5y}\color{green}{-9} =\color{blue}{-y}\color{green}{+1}\)

OpenStudy (anonymous):

Ok so now it's -4y-9=1?

OpenStudy (studygurl14):

yep. now combine the -9 a nd 1

OpenStudy (anonymous):

-4y=-8

OpenStudy (studygurl14):

not quite. Do 1 + 9

OpenStudy (anonymous):

oh -4y-8

OpenStudy (anonymous):

Do I divide them?